Monochromatic radiation of wavelength 'x' is incident on a hydrogen sample in ground state . Hydrogen atoms absorb a fraction of light and subsequently emit radiation of six wavelengths. Find the value of 'x'????

Grade:12th pass

1 Answers

Arun
5 years ago
Dear student
 
 
from the eqn. for ground state
n(n-1)/2 = 6
n^2 – n – 12 = 0
n^2 – 4n + 3n – 12 = 0
(n-4) (n+3) = 0
n = 4
Therfore,
1/x = R*1^2(1/1^2 – 1/4^2)
1/x = R*(1 – 1/16) => 15R/16
x = 16/15R   = 16*10^7/(15 *1.09) is the answer.    = 0.978593 * 10^7 meter.
 
Regards
Arun (askIITians forum expert)